OpenStudy (anonymous):

WILL AWARD MEDALS Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6

OpenStudy (anonymous):

@gurrilabill @Shay17 Can you guys help me out?

OpenStudy (anonymous):

input (k+1) where the n is and it will help you figure it out i had that same question :)

OpenStudy (amistre64):

first, what is your basis step?

OpenStudy (amistre64):

... is true for all positive integers n, or show why it is false. since positive integers are: 1,2,3,4,5,.... is this true for n=1?

OpenStudy (amistre64):

try it out: 4n( 4n + 2) = 4(4n+1)(8n+7)/6 4(6) = 4(5)(15)/6 6 = 5(15)/6 36 = 5(15) 36 = 50+25, is that a true statement?

OpenStudy (anonymous):

It's false.

OpenStudy (anonymous):

So my answer is false? Or do I need to take further steps? @amistre64

OpenStudy (amistre64):

... is true for all positive integers n, or show why it is false. since it is not true for at least one positive integer (n=1), then we have proven it to be false.

OpenStudy (amistre64):

there is no need to continue to try proving it true when it is false from the start .... wouldnt you agree?

OpenStudy (anonymous):

Yes, I agree (: Thank you so much for your help!